# SlipStream: Adapting Pipelines for Distributed Training of Large DNNs Amid Failures

### v1

@@ 3.1. Adaptive Pipelining: Working Around Failures
The scheduling challenge: In theory, the execution schedule of the two peer workers has enough bubbles to do the computations for failed node $W_{1\_2}$ as there are 18 idle slots in total in Figure 3(a). However, these bubbles are concentrated in the warm-up and cool-down phase of the 1F1B schedule. Adaptive Pipelining needs to reroute micro-batches to peer workers $W_{0\_2}$ and $W_{2\_2}$ mostly in the middle of their steady-state schedule, which is optimized to be bubble-free for efficiency (Narayanan et al., 2019). As shown in Figure 3(b), this has a significant performance impact. The additional micro-batches to peer workers and the dependencies between forward and backward pass computations lead to 36 time steps per iteration. In other words, the failure of 8.3% of workers (1 out of 12) leads to a 33% slowdown in iteration time.
3.2. Decoupled BackProp: Filling Unused Bubbles
@@
Figure 6. The Staggered Optimizer allows SlipStream to optimize pipeline schedule across training iterations.
@@
Because $B_{Weight}$ is dependence-free, it can largely be deferred until the end of the training iteration. Hence, we can take advantage of idle slots in the cool-down phase of the 1F1B schedule for $B_{Weight}$ computations, freeing slots in the steady stage for the re-routed work of the failed worker. This enables SlipStream to absorb the additional computational demands imposed by failures with minimal overheads.
The memory challenge. Unfortunately, Decoupled BackProp increases memory pressure. Decoupling the computation of $B_{Weight}$ requires that intermediate data is stored for potentially extended periods on each worker. For instance, as shown in Figure 5, decoupling the backward pass for micro batch 1 necessitates retaining its intermediate data until time step 25. To avoid memory exhaustion, SlipStream applies Decoupled BackProp selectively only when it can mitigate the overheads of Adaptive Pipelining. We also capitalize on the observation that memory imbalance exists among the pipeline stages (Kim et al., 2023). To ensure that all pipeline stages are fully utilized, earlier stages need to allocate additional memory to process more forward micro-batches than later stages (Narayanan et al., 2019). By exploiting this available surplus memory, SlipStream can effectively offset some of the memory demands of Decoupled BackProp without incurring additional costs.
@@ 3.3. Staggered Optimizer: Accessing More Bubbles
Decoupled BackProp makes efficient use of the cool-down bubbles. However, as seen in Figure 6, the warm-up bubbles remain underutilized due to their placement after the synchronous optimizer step of the previous iteration. To exploit warm-up bubbles, we make the critical observation that optimizer steps for different pipeline stages are independent of each other. The Staggered Optimizer thus shifts the timing of the optimizer step across pipeline stages to better utilize the bubbles that occur during the warm-up phase of the subsequent training iteration. Effectively, this staggering provides SlipStream more bubbles in the cool-down phase to hide the overhead of compensating for failed workers.
@@
Put together, these three mechanisms allow SlipStream to flexibly optimize the use of computational resources across pipelines, ensuring that each stage operates near its capacity and maintains consistent utilization throughout training.
3.4. Supporting Multiple Failures
SlipStream’s techniques support multiple failures. SlipStream guarantees continued training up to $DP-1$ simultaneous worker failures, where $DP$ is the number of data-parallel pipelines, as one functional worker exists for each pipeline stage. SlipStream can probabilistically sustain more failures, as long as there is at least one functional worker per stage across all data parallel pipelines. Figure 7(b) shows how SlipStream can recover from more than $DP-1$ failures. Despite 8 simultaneous failures – 2/3 of the GPUs! – SlipStream continues training.
@@ 3.4. Supporting Multiple Failures
(b) SlipStream can continue training if a functional worker exists per stage.
Figure 7. Example of SlipStream’s fault tolerance guarantees.
As workers are repaired, SlipStream can re-insert them back into the adapted pipeline schedule, reducing the probability of high counts of concurrent failures. In the rare event that simultaneous failures affect a single data parallel group as shown in Figure 7(a), SlipStream resorts to current resilience solutions: calculate an efficient hybrid-parallel scheme for the remaining nodes, configure it across all workers by restoring from a recent checkpoint, and resume training at the speed the new parallelism scheme allows.
Multi-GPU Servers and Tensor Parallelism. Training systems commonly employ servers with multiple GPUs. A favoured configuration involves HGX/DGX servers with 8 GPUs connected in an all-to-all manner using NVLink and NVSwitches (DGX, 2024). Similar to Oobleck and Bamboo, we consider a whole server as the unit of failure. This approach is pragmatic because servicing even one faulty GPU board necessitates taking the entire server offline. Moreover, software faults, CPU issues, connectivity problems, and power or cooling issues can also incapacitate a whole server. We also need to account for the level of tensor parallelism. Most training systems apply tensor parallelism within a multi-GPU server to leverage the high-bandwidth NVLink connections. In cases where tensor parallelism spans multiple servers, SlipStream, along with Oobleck and Bamboo, identifies the collective group of tensor-parallel servers as a unit of failure.
4. SlipStream Design
@@ 4.1. System Overview
SlipStream consists of three key components, the Profiler, the Executors, and critically the Planner.
@@
Profiler. When a large training job is first submitted, SlipStream runs a short profiling job to collect key performance statistics such as the average micro-batch latency for the forward and backward passes, memory requirements for activations and gradients, and inter-node communication bandwidth. The profiling job executes a small number of training iterations, 100 by default, and typically takes a few minutes. These statistics are used by the Planner.
@@
Planner. The Planner generates near-optimal schedules for the training job under various failure scenarios. It uses dynamic programming (DP) and mixed integer linear program (MILP) formulations that implement the SlipStream techniques (Section 3). We generate a pipeline schedule for each number of tolerated simultaneous failures. Each plan is agnostic to which specific worker(s) fail. It encodes the specific pipeline schedule including micro-batch assignment, the order of forward and backward tasks (e.g., ForwardPass, BackwardWeightPass), and the communication plan (e.g., RecvActiv, SendGrad, etc.). The plans are stored in distributed fault-tolerant storage (e.g., etcd) to be used by runtime Executors.
@@
By default, we generate up to $DP-1$ plans, which accounts for the $DP-1$ simultaneous failures that SlipStream can always handle. Since SlipStream can probabilistically sustain many more failures, we can also generate plans for up to a user-defined fault tolerance threshold. The Planner runs relatively quickly, requiring 53 minutes to generate plans for up to 512 concurrent failures for a cluster of 2048 GPUs. Nevertheless, the training job can begin training in parallel as most plans will not be needed immediately.
@@
Executors. SlipStream operates with an Executor on each GPU node, tasked with implementing the training plan specific to that node. Overseeing these Executors, a centralized Coordinator orchestrates the overall training operation by distributing the appropriate execution plans to each Executor. Additionally, the Coordinator actively monitors the status of all Executors to ensure their continuous operation. In the event of a failure, the Coordinator reassigns the tasks impacted by failure to the most suitable location, as advised by the Planner (refer to Section 4.2.1), and then directs the remaining Executors to resume training according to the updated plan that now reflects the updated count of non-operational workers. Training resumes from the iteration during which the failure was identified. If failures are not promptly detected, restoring a checkpoint from remote storage may become necessary. To enhance reliability, the Coordinator itself is safeguarded against failures through active replication.
4.2. SlipStream Planner
@@
Given the cluster configuration, training job, profiling statistics, and set of failed workers, the Planner calculates a valid pipeline schedule that minimizes the training iteration latency using the fault-free workers. The Planner uses two phases, Normalization followed by Solution.
@@
SlipStream generates an execution plan for a given number of failures, regardless of which worker(s) fails, as solving ILP problems for the combinatorial number of possible failure locations is intractable. The first phase normalizes the specific failures by migrating failed nodes to a pre-determined and ideal set of locations within the pipeline schedule. For example, if any single node fails in the example in Figure 7, the Planner will always swap the node with an ideal location, e.g. $W_{2\_3}$. The Planner generates a mapping of the ideal migration locations for each possible number of simultaneous failures. Unlike current solutions which require a complete reconfiguration of a data-parallel pipeline, normalization only requires a point-to-point copy of model parameters to swap the location of two workers in the pipeline for each failure. The Solution Phase leverages an MILP to derive an optimized schedule for each normalized case.
4.2.1. Normalization Phase
@@
The intuition behind the Normalization Phase is to a) distribute failures across different peer groups (to maximize fault tolerance and load balance extra work across peer groups) and b) move failures to later pipeline stages with more bubbles (to maximize the opportunity to hide overheads).
Normalization relies on dynamic programming (DP) to determine the ideal pipeline location to migrate faults for given a number of failures. We use a DP algorithm as it can greedily determine the ideal migration strategy with very low runtime ($O(NF^{2})$ complexity for all solutions up to $F$ failures with $N$ stages per pipeline). The DP algorithm is based on a model that calculates the cost in terms of the extra bubbles used to accommodate re-routed micro-batches.
DP Formulation
Algorithm 1 Normalization Phase DP Algorithm
1:$N\leftarrow$ Number of pipeline stages.
2:$F\leftarrow$ Total number of failures.
3:$C\leftarrow$ an Nx(F+1) array $\triangleright$ Costs
4:$A\leftarrow$ an Nx(F+1) array $\triangleright$ Assignments
5:for $i\in\{0,\dots,N-1\}$ do
6:     for $f\in\{0,\dots,F\}$ do
7:         if $i==0$ then
8:              $C[i][f]=cost(i,f)$
9:              $A[i][f]=[f]$
10:         else
11:              $C[i][f]=\min_{x\leq f}(C[i-1][f-x]+cost(i,x))$
12:              $x=\operatorname*{arg\,min}_{x\leq f}(C[i-1][f-x]+cost(i,x))$
13:              $A[i][f]=concat(A[i-1][f-x],x)$
14:         end if
15:     end for
16:end for
17:$R=A[N-1][F]$
18:return $R$
@@
Given the total number of failures to handle, $F$, and the total number of pipeline stages within a data parallel pipeline, $N$, the DP algorithm returns a list $R$ of length $N$, such that $\Sigma_{i}R[i]=F$. Each index $R[i]$ specifies the ideal number of failures that is assigned to pipeline stage $i$ across all data parallel pipelines; the specific pipeline assignment does not affect performance and can be arbitrary. For example, if $R[3]=2$ for the example in Figure 7(a), the ideal migration would be to swap the failed nodes with any two nodes in stage $3$, for example $W_{0\_3}$ and $W_{2\_3}$.
For each DP step, $C[i][f]$ represents the total cost, in terms of bubbles used, to handle $f$ failures at all stages between 00 and $i$. Meanwhile, each $A[i][f]$ contains a list that denotes the portion of the $f$ failures assigned to each stage from 00 to $i$, as described for $R$ – in other words $len(A[i][f])==i+1$. Each DP step uses a recurrence relation that aims to minimize the cost of handling $f$ failures within the first $i$ pipeline stages across all of the data parallel pipelines, while $A[i][f]$ denotes the assignment of failures that achieves this cost. The ideal assignment to handle $F$ failures across all workers is simply $R=A[N-1][F]$.
To estimate the cost, $cost(i,x)$ is a function that returns the extra bubbles used to handle the extra micro-batches from $x$ failures in the $i$-th stage (i.e., peer group). While we can use the MILP described in Section 4.2.2 to estimate the cost, the Planner can also rely on a heuristic to estimate the cost in order to avoid longer solve times (as Algorithm 1 requires $O(NF)$ calls to the cost function). The heuristic captures the benefit of the techniques described in Section 3. It creates a pipeline schedule that is similar to 1F1B, except that the heuristic’s 1F1B backward pass only schedules $B_{input}$. Then, when there exists a gap (both in terms of time and memory) that is able to fit $B_{weight}$ (or the $F$ and $B_{input}$ stages are exhausted), $B_{weight}$ operations are scheduled opportunistically.
4.2.2. Solution Phase
@@
Next, the Planner uses the normalized failure locations and the profiled statistics as inputs to an MILP which performs schedule optimization. At a high level, the MILP determines how to schedule re-routes micro-batches across peers of a failed worker alongside regular micro-batches in those peers (Adaptive Pipelining). The formulation accounts for the latency for communicating gradients across workers, task dependencies in forward and backward passes, and the memory usage in each worker. The definition of task dependencies also allows the solution to leverage Decoupled BackProp and Staggered Optimizer techniques.
MILP Formulation
We aim to minimize the makespan of a single training iteration which is repeated throughout the training process, while the constraints are task dependencies and the memory capacity of workers.
@@
Notation. An operation within a training iteration can be explicitly identified across workers using a 5-tuple $(i,j,k,c,k_{s})$. $i$ represents the pipeline stage within each data parallel pipeline, while $k$ represents the operation’s original data parallel pipeline (prior to failures). $j$ is the micro-batch ID in the training iteration. $c$ identifies the different operations for each micro-batch, $c\in\{F,B_{input},B_{weight}\}$. Finally, $k_{s}$ denotes the peer pipeline in the event that a micro-batch originally intended for $k$ runs on $k_{s}$ (note that $k$ can equal $k_{s}$). For example, a micro-batch originally intended to execute on $W_{2\_3}$, but was re-routed to a peer $W_{1\_3}$ in Figure 7, would have $i=3$, $k=2$, and $k_{s}=1$. Thus, $(i,k_{s})$ uniquely identifies the worker for a given operation.
Inputs. The MILP takes as input the micro-batch to worker assignments in an iteration. Specifically, the Planner takes the normalized failure locations and assigns all micro-batches corresponding to failed workers evenly across peer workers within the same peer group, alongside the original (non-peer) micro-batches. Thus, the Planner produces a binary mapping $S_{i,j,k}^{k_{s}}\in\{0,1\}$, which indicates if a micro-batch $(i,j,k)$ should be executed on pipeline $k_{s}$. Note that $S_{i,j,k}^{k}=1$ if the micro-batch executes on the original pipeline. We ensure that each operation is assigned only one pipeline, i.e., $\sum_{\begin{subarray}{c}k_{s}\end{subarray}}S_{i,j,k}^{k_{s}}=1$.
@@
The MILP also takes as input profiled statistics. First, $T_{c}$ represents the computational time for each operation $c$. $T_{comm}$ represents the communication latency to send activations or gradients across workers [^1] We use a single value as activations and gradients are the same size.. Next, the MILP requires the change in memory utilization on worker $(i,k,k_{s})$ caused by executing an operation $(i,j,k,c,k_{s})$ as $\Delta M_{i,j,k,c}^{k_{s}}$.
@@
Here, $A_{B}$ is the profiled size of activation stored at the end of forward pass. $A_{B_{input}}$ and $A_{B_{Weight}}$ is size of the activation stored at the end of backward-input and backward-weight pass, respectively. Upon completion of backward-input pass and backward-weight pass, we free activation $A_{B_{Input}}$ and $A_{B_{Weight}}$ respectively.
Variables. The goal of the MILP is to identify the exact ordering of operations, which can be used to generate a pipeline schedule. To do this, we introduce an ordering variable $O_{(i,j,k,c,k_{s})\rightarrow(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})}\in\{0,1\}$ between each pair of operations $(i,j,k,c,k_{s})$ and $(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})$. Each variable is 1 if $(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})$ is scheduled any time after $(i,j,k,c,k_{s})$ and 0 otherwise. Ordering is only required within a stage, and between computation phases of same micro-batch (e.g., between the $F$ and $B_{input}$ pass). Finally, we define the set of variables $E_{i,j,k,c}^{k_{s}}$ as the ending time of operation $(i,j,k,c,k_{s})$.
Objective. The goal of the MILP is to minimize the makespan of a single training iteration (which is repeated throughout the training process). This can be done by determining the sequence and time of operations through the training pipeline via the sets of variables $O$ and $E$. This objective can thus be formulated as:
@@
Finally Equation 6 calculates the activation memory required at any given time on worker $(i,k_{s}^{\prime})$, and constrains operations such that the total memory is below an $M_{limit}$ to avoid OOMs.
4.3. Implementation

### v2

@@ 3.2. Decoupled BackProp: Filling Unused Bubbles
Figure 6. The Staggered Optimizer allows ReCycle to optimize pipeline schedule across training iterations.
@@
Because $B_{Weight}$ is dependence-free, it can largely be deferred until the end of the training iteration. Hence, we can take advantage of idle slots in the cool-down phase of the 1F1B schedule for $B_{Weight}$ computations, freeing slots in the steady stage for the re-routed work of the failed worker. This enables ReCycle to absorb the additional computational demands imposed by failures with minimal overheads.
The memory challenge. Unfortunately, Decoupled BackProp increases memory pressure. Decoupling the computation of $B_{Weight}$ requires that intermediate data is stored for potentially extended periods on each worker. For instance, as shown in Figure 5, decoupling the backward pass for micro batch 1 necessitates retaining its intermediate data until time step 25. To avoid memory exhaustion, ReCycle applies Decoupled BackProp selectively only when it can mitigate the overheads of Adaptive Pipelining. We also capitalize on the observation that memory imbalance exists among the pipeline stages (Kim et al., 2023). To ensure that all pipeline stages are fully utilized, earlier stages need to allocate additional memory to process more forward micro-batches than later stages (Narayanan et al., 2019). By exploiting this available surplus memory, ReCycle can effectively offset some of the memory demands of Decoupled BackProp without incurring additional costs.
@@ 3.3. Staggered Optimizer: Accessing More Bubbles
Decoupled BackProp makes efficient use of the cool-down bubbles. However, as seen in Figure 6, the warm-up bubbles remain underutilized due to their placement after the synchronous optimizer step of the previous iteration. To exploit warm-up bubbles, we make the critical observation that optimizer steps for different pipeline stages are independent of each other. The Staggered Optimizer thus shifts the timing of the optimizer step across pipeline stages to better utilize the bubbles that occur during the warm-up phase of the subsequent training iteration. Effectively, this staggering provides ReCycle more bubbles in the cool-down phase to hide the overhead of compensating for failed workers.
@@
Put together, these three mechanisms allow ReCycle to flexibly optimize the use of computational resources across pipelines, ensuring that each stage operates near its capacity and maintains consistent utilization throughout training.
3.4. Supporting Multiple Failures and Re-Joins
@@
(b) ReCycle can continue training if a functional worker exists per stage.
Figure 7. Example of ReCycle’s fault tolerance guarantees.
ReCycle supports multiple failures and rejoins. ReCycle guarantees continued training up to $DP-1$ simultaneous worker failures, as in the worst case $DP$ failures can disable all the data-parallel peers of a pipeline stage. However, ReCycle can probabilistically sustain more failures, as long as there is at least one functional worker per stage across all data parallel pipelines. Figure 7(b) shows how ReCycle can recover from more than $DP-1$ failures. Despite 8 simultaneous failures – 2/3 of the GPUs! – ReCycle ensures continual training.
As workers are repaired, ReCycle can re-insert them back into the adapted pipeline schedule, reducing the probability of high counts of concurrent failures. Unlike unexpected failures, planned worker additions occur at iteration boundaries to overlap any data copying overhead with the previous iteration. Once a worker rejoins, ReCycle ceases rerouting micro-batches and sends all micro-batches from the previous stage to the recovered worker.
In the worst-case scenario, where simultaneous failures impact a single data-parallel group (as shown in Figure 7(a)), ReCycle falls back on existing resilience strategies: it calculates an efficient hybrid-parallel scheme for the remaining nodes, configures it across all workers by restoring from a recent checkpoint, and resumes training at the speed supported by the new parallelism.
Multi-GPU Servers and Tensor Parallelism. Training systems commonly employ servers equipped with multiple GPUs. A favoured configuration features HGX/DGX servers with 8 GPUs connected in an all-to-all manner using NVLink and NVSwitches (DGX, 2024). Similar to Oobleck and Bamboo, we treat an entire server as the unit of failure. This approach is pragmatic because servicing even one faulty GPU board necessitates taking the entire server offline. Additionally, issues such as software malfunctions, CPU issues, connectivity disruptions, and power or cooling issues can incapacitate an entire server. Most training systems implement tensor parallelism within a multi-GPU server to leverage the high-bandwidth NVLink connections. However, a tensor parallel group can extend across multiple servers. In such cases, ReCycle, along with Oobleck and Bamboo, classifies the entire group of tensor-parallel servers as a single unit of failure.
4. ReCycle Design
@@ 4.1. System Overview
ReCycle consists of three key components, the Profiler, the Executors, and critically the Planner, as shown in Figure 8.
@@
Profiler. When a large training job is first submitted, ReCycle runs a short profiling job to collect key performance statistics such as the average micro-batch latency for the forward and backward passes, memory requirements for activations and gradients, and inter-node communication bandwidth. The profiling job executes a small number of training iterations, 100 by default, and typically takes a few minutes. These statistics are used by the Planner.
@@
Planner. The Planner generates pipeline schedules for the training job under various failure scenarios. It uses dynamic programming and mixed integer linear programming (MILP) to implement the ReCycle techniques (Section 3). We generate a pipeline schedule for each number of tolerated simultaneous failures. Each plan is agnostic to which specific worker(s) fail. It encodes the specific pipeline schedule, specifying micro-batch assignments, the sequence of forward and backward tasks, and communication of activations and gradients. The plans are stored in distributed fault-tolerant storage (e.g., etcd) to be used by runtime Executors.
@@
By default, we generate up to $DP-1$ plans, which accounts for the $DP-1$ simultaneous failures that ReCycle can always handle. Since ReCycle can probabilistically sustain many more failures, we can also generate plans for up to a user-defined fault tolerance threshold.
@@
Executors. ReCycle operates with an Executor on each GPU node, tasked with implementing the training plan specific to that node. Overseeing these Executors, a centralized Coordinator orchestrates the overall training operation by distributing the appropriate execution plans to each Executor. Additionally, the Coordinator actively monitors the status of all Executors to ensure their continuous operation. In the event of a failure, the Coordinator reassigns the tasks impacted by failure to the most suitable location, as advised by the Planner (refer to Section 4.2.1), and then directs the remaining Executors to resume training according to the updated plan that now reflects the updated count of non-operational workers. Training resumes from the iteration during which the failure was identified. If failures are not promptly detected, restoring a checkpoint from remote storage may become necessary. To enhance reliability, the Coordinator itself is safeguarded against failures through active replication.
4.2. Planner
@@
Given the cluster configuration, training job, profiling statistics, and failed workers, the Planner calculates an adaptive schedule that minimizes the training iteration latency using the fault-free workers. It does this in two phases, Failure Normalization followed by Adaptive Schedule Generation.
@@
To avoid solving an MILP for the combinatorial number of possible failure locations, the Failure Normalization phase first normalizes each given number of failures by calculating the suitable location within the pipeline schedule to migrate each failure. For example, if any single node fails in the example in Figure 7, the Planner will swap the node with the calculated ideal location, e.g., $W_{2\_3}$. Unlike current solutions which require a complete reconfiguration of a data-parallel pipeline (Jang et al., 2023), normalization only requires a point-to-point copy of model parameters to swap the location of two workers in the pipeline for each failure. The Adaptive Schedule Generation then leverages a MILP to derive an adaptive pipeline schedule for each normalized case.
4.2.1. Failure Normalization
Algorithm 1 Failure Normalization
1:$DP\leftarrow$ Number of data-parallel pipelines.
2:$PP\leftarrow$ Number of pipeline stages.
3:$MB\leftarrow$ Number of microbatches per pipeline.
4:$F\leftarrow$ Total number of failures.
5:$O\leftarrow$ an PP $\times$ (F+1) array $\triangleright$ Rerouting Overheads
6:$A\leftarrow$ an PP $\times$ (F+1) array $\triangleright$ Assignments
7:
8:procedure failure_reordering($DP,PP,MB,F$)
9:     for $i\in\{0,\dots,PP-1\}$ do
10:         for $f\in\{0,\dots,F\}$ do
11:              if $i==0$ then
12:                  $O[i][f]=\textsc{COST}(f)$
13:                  $A[i][f]=[f]$
14:              else
15:                  $x=\operatorname*{arg\,min}_{x\leq f}\Bigl{(}O[i-1][f-x]$
[…]
17:                  $O[i][f]=O[i-1][f-x]+\textsc{COST}(x)$
18:                  $A[i][f]=concat\Bigl{(}A[i-1][f-x],x\Bigr{)}$
19:              end if
20:         end for
21:     end for
22:return $A[PP-1][F]$
23:end procedure
24:
25:procedure cost($f$)
26:     if $f>0$ then
27:         return $\min\Bigl{(}0,MB\times f\times 3-(DP-f)\times(PP-1)\times 3\Bigr{)}$
28:     end if
29:     return 00
30:end procedure
@@ 4.2.1. Failure Normalization
The intuition behind Failure Normalization, shown in Algorithm 1, is twofold: a) distribute failures across different peer groups to enhance fault tolerance and evenly balance the additional workload, and b) shift failures to later pipeline stages with more bubbles to maximize the opportunity to hide re-routing overheads.
@@
Failure Normalization uses dynamic programming to compute a migration strategy for handling a given number of failures. Given a total of $F$ failures and $PP$ pipeline stages, it returns a list $A$ of length $PP$, where the sum of elements in $A$ equals $F$. Each $A[i]$ specifies the number of failures assigned to pipeline stage $i$ across all data parallel pipelines, with the specific pipeline assignments being arbitrary and not impacting performance. For example, if $A[3]=2$, two failures would be assigned to stage 3, and we could swap the failed nodes with any two from that stage, such as $W_{0\_3}$ and $W_{2\_3}$ in Figure 7(b).
The overhead of handling $f$ failures at stages from 00 to $i$ is represented by $O[i][f]$, while $A[i][f]$ indicates how the $f$ failures are distributed across these stages, with $A[i][f]$ being of length $i+1$. The recurrence relation minimizes overhead by finding the optimal failure assignment for $f$ failures across the first $i$ stages, and the best assignment for $F$ failures is given by $A[PP-1][F]$.
To compute $COST(i,x)$, we estimate the additional time slots needed for handling extra micro-batches due to $x$ failures in the $i$-th stage using the heuristic from line 27. While the MILP in Section 4.2.2 could be used, Failure Normalization opts for the heuristic to reduce computation time. The complexity of determining the migration strategy for $F$ failures and $PP$ pipeline stages is $O(PP\times F)$.
4.2.2. Adaptive Schedule Generation
@@
Next, the Planner uses normalized failure locations and profiled statistics as inputs for an MILP to generate adaptive schedule. At a high level, the MILP determines how to re-route micro-batches across peers of a failed worker, integrating both regular and re-routed micro-batches (Adaptive Pipelining). It considers communication latency $T_{comm}$, micro-batch latency for forward, backward input and weight pass – $T_{F}$, $T_{B_{Input}}$ and $T_{B_{Weight}}$ respectively, task dependencies in forward and backward passes, and memory usage, leveraging the Decoupled BackProp and Staggered Optimizer techniques.
@@
Notation. Each operation in a training iteration is denoted by the 5-tuple $(i,j,k,c,k_{s})$. Here, $i$ represents the pipeline stage within each data parallel pipeline, and $k$ indicates the operation’s original data parallel pipeline before any failures. The variable $j$ denotes the micro-batch ID in the training iteration, and $c$ specifies the type of operation for each micro-batch, where $c\in\{F,B_{input},B_{weight}\}$. Finally, $k_{s}$ specifies the peer pipeline that executes a micro-batch originally intended for $k$; which can be the same as $k$. For instance, a micro-batch ID $14$ originally scheduled for $W_{2\_3}$ but rerouted to peer $W_{1\_3}$ would be identified with $i=3$, $j=14$, $k=2$, and $k_{s}=1$.
Inputs. We use micro-batch assignments for each worker as inputs. The Planner reassigns micro-batches from failed workers to peer workers within the same group while retaining the original micro-batches. This results in a binary mapping $S_{i,j,k}^{k_{s}}\in\{0,1\}$, indicating whether a micro-batch $(i,j,k)$ should run on pipeline $k_{s}$. Each operation is assigned to exactly one pipeline: $\sum_{k_{s}}S_{i,j,k}^{k_{s}}=1$.
@@
Additionally, we use profiled statistics, which include $T_{c}$ for the computational time of each operation $c$, $T_{comm}$ for communication latency of activations or gradients [^1] We use a single value as activations and gradients are the same size., and $\Delta M_{i,j,k,c}^{k_{s}}$ for the change in memory utilization on worker $(i,k,k_{s})$ due to the execution of operation $(i,j,k,c,k_{s})$.
@@
Here, $A_{B}$ is the profiled size of activation at the end of the forward pass. $A_{B_{input}}$ and $A_{B_{weight}}$ are the sizes of gradients at the end of the backward-input and backward-weight passes, respectively. We free $A_{B_{input}}$ and $A_{B_{weight}}$ after completing their respective backward passes.
Variables. We define a binary variable $O_{(i,j,k,c,k_{s})\rightarrow(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})}\in\{0,1\}$ to represent ordering between pair of operations $(i,j,k,c,k_{s})$ and $(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})$. This variable is 1 if $(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})$ is scheduled after $(i,j,k,c,k_{s})$, and 0 otherwise. Ordering is needed only within a stage and between computation phases of the same micro-batch (e.g., $F$ and $B_{input}$). Additionally, $E_{i,j,k,c}^{k_{s}}$ represents the ending time of operation $(i,j,k,c,k_{s})$.
Objective. Our objective is to minimize the makespan of a single training iteration while adhering to task dependencies and memory constraints. This involves determining the sequence and timing of operations throughout the training pipeline using the sets of variables $O$ and $E$. Thus, the objective can be formulated as follows:
@@
Finally, Equation 6 calculates the activation memory required at any given time on worker $(i,k_{s}^{\prime})$, and constrains operations such that the total memory is below an $M_{limit}$.
5. Implementation